Diego Company manufactures one product that is sold for $77 per unit in two geographic regions—the East and West regions. The following information pertains to the company’s first year of operations in which it produced 48,000 units and sold 43,000 units.

Variable costs per unit:
Manufacturing:
Direct materials $ 27
Direct labor $ 12
Variable manufacturing overhead $ 3
Variable selling and administrative $ 5
Fixed costs per year:
Fixed manufacturing overhead $ 864,000
Fixed selling and administrative expense $ 456,000

The company sold 33,000 units in the East region and 10,000 units in the West region. It determined that $220,000 of its fixed selling and administrative expense is traceable to the West region, $170,000 is traceable to the East region, and the remaining $66,000 is a common fixed expense. The company will continue to incur the total amount of its fixed manufacturing overhead costs as long as it continues to produce any amount of its only product.

14. Diego is considering eliminating the West region because an internally generated report suggests the region’s total gross margin in the first year of operations was $50,000 less than its traceable fixed selling and administrative expenses. Diego believes that if it drops the West region, the East region's sales will grow by 5% in Year 2. Using the contribution approach for analyzing segment profitability and assuming all else remains constant in Year 2, what would be the profit impact of dropping the West region in Year 2?

15. Assume the West region invests $38,000 in a new advertising campaign in Year 2 that increases its unit sales by 20%. If all else remains constant, what would be the profit impact of pursuing the advertising campaign?
Answer #1
Unit selling price 77
Less:Unit variable cost 47 =27+12+3+5
Unit Contribution margin 30
14
Loss in Contribution margin of West region -300000 =-10000*30
Increase in Contribution margin of West region 49500 =33000*5%*30
Avoidable fixed selling and administrative expense 220000
Net change in profits -30500
Profits will decrease by 30500
15
Incremental contribution margin 60000 =10000*20%*30
Less: Advertising costs -38000
Net change in profits 22000
Profits will increase by 22000